Shift report 20000719 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Wed Jul 19 2000 - 16:00:08 EDT

  • Next message: http account: "Shift report 20000720 16:00-24:00"

    ------------------------------------------------------------
    Date: Wed Jul 19 16:00:07 2000, Shift: 08:00-16:00
    Supervisor: YKL, with: YB
    Log book#3 pages 29 - 30
    
    Shift Summary:
    ------------------------------------------------------------
    RHIC had quench interlock problems and power supply problems 
    all morning. RHIC had injection studies for the balance 
    of the shift. There was no success in ramping and storing
    beams.
    
    pii4 is rebooted. pii3 mulfuctions and Betty worked on it.
    
    HV for MULT tile01 tripped. When it was turned on the
    measured voltage is a few volts off. 
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Wed Jul 19 2000 - 16:03:23 EDT